If M,N are R-module then M x N is also R-module
1
Expert's answer
2020-11-03T17:07:23-0500

M x N the following operations are set:

"(m_1,n_1)+(m_2,n_2)=(m_1+m_2,n_1+n_2)\\
r*(m_1,n_1)=(r*m_1,r*n_1)\\
r\in R,\ m_1,m_2\in M,n_1,n_2\in N"

Since operations on M x N are "distributed" into two parts in M and N separately, it is quite obvious that M x N is an R-module. For example, let's check the commutativity:

"(m_1,n_1)+(m_2,n_2)=(m_1+m_2,n_1+n_2)=\\
=(m_2+m_1,n_2+n_1)=(m_2,n_2)+(m_1,n_1)"